• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name mcnulty

    Log entry time 09:16:01 on September16,2009

    Entry number 289894

    Followups:

    keyword=Work done during access/ spot++ problem investigation

    I'm out of the hall. I could not find any problem with the Larm DAQ. I did a cosmic run with s0, hapdet, and VDCs on (24726) and as far as I can tell everything is normal. However, spot++ (and spot_L) for the Left is still giving the same errors.

    I also took a clock trigger run in the Rarm (4602) and found spot_R to work as before -- with the strange/missing results for BPMA data and normal/good results for BPMB.

    In the right arm I did not find that the BPM signal cable was un-seated or sagging at the ADC input--which was suspected as causing the strange spot_R results. However, the raster current distributions from spot_R look normal/good and it seems we can use these to center the raster on the holey Carbon target and goto production.

    Bob is currently verifying/checking that the Larm daq is functioning properly. As soon as we are confident, we will goto beam permit.

    We are currently taking cosmic runs in both arms 24729/4603 and the data looks good to me.

    Calling MCC for Beam Permit. Will use spot_R for raster verification.